OpenStudy (anonymous):

Find the derivative of f(x) = sin(x) at x=3.142/4 using numerical derivative. Also calculate the errors. Use h=0.5, 0.25, 0.125, 0.0625. (If possible, can u show me the way how to calculate more detail.) thank you.

OpenStudy (anonymous):

Derivative of f(x)=sin(x) is f'(x)=cos(x). At pi/4, cos(x)=sqrt(2)/2, which should be roughly .707. Sounds like it wants you to use the limit definition though if it's providing you with values for h. So you have to do f(x+h)-f(x)/h for each h, presumably.